How many grams of F are in 195 g CaF2?
1
Expert's answer
2018-03-16T09:14:41-0400
Answer:
n ( CaF2) = m/M
n (CaF2) = 195/78 = 2.5 mole
n (F) = 5 mole
m (F) = 5*19 = 95 g